OpenStudy (anonymous):

A box of groceries requires 5.0 newtons of force to lift it up 1.0 meter. How much work is done? 0.0 J 1.0 J 2.0 J 5.0 J

hartnn (hartnn):

Work done = Force times Displacement W = F*d here, F = 5 , d = 1 just plug in values!

hartnn (hartnn):

you will directly get answer in Joules , J :)

OpenStudy (anonymous):

I cant get the answer

hartnn (hartnn):

W = F*d = 5*1 = .... ?

OpenStudy (anonymous):

so the answer is number 2

hartnn (hartnn):

whats 5*1 ?

OpenStudy (anonymous):

1

hartnn (hartnn):

how ?

OpenStudy (anonymous):

what is 5*1 ? i dont understand

OpenStudy (anonymous):

are you trying to say 5x1?

hartnn (hartnn):

yes

OpenStudy (anonymous):

kay so 5x1=1

OpenStudy (anonymous):

so is one that answer

hartnn (hartnn):

how did you know 5x1 = 1 ?

OpenStudy (anonymous):

lol i mean 5

OpenStudy (anonymous):

so the answer is 5

hartnn (hartnn):

yes

OpenStudy (anonymous):

thank you ill give you a medal

hartnn (hartnn):

you're welcome :)